A curling stone is given an initial velocity of 18.4m/s across ice rink. If the coefficient of friction between the stone and the ice s 0.054, how far will the stone travel before it comes to a stop.

1
Expert's answer
2022-03-01T17:33:08-0500

The distance can be found as follows:


"d = \dfrac{W}{F}"

where "W" is the work done by friction force "F". According to the energy-work theorem, the work is equal to the change in kinetic energy "W = \Delta K". Since the stone starts from "v = 18.4m\/s" and comes to the rest, the change in kinetic energy is the following:


"\Delta K = \dfrac{mv^2}{2}"

where "m" is the mass of the stone.

The friction force is given as follows:


"F = \mu m g"

where "\mu=0.054,g = 9.8m/s^2". Thus, find:


"d = \dfrac{mv^2}{2\mu mg} = \dfrac{v^2}{2\mu g} = \dfrac{18.4^2}{2\cdot 0.054\cdot 9.8} \approx 320m"

Answer. 320m.